OpenStudy (anonymous):

a straight line has the equation x-2y+4=0 what angle does the line make with x+3y=0

OpenStudy (amistre64):

if you can compare slope, that would be useful to a solution

OpenStudy (amistre64):

 \[cos(a)=\frac{a.b}{|a||b|}\]

OpenStudy (amistre64):

prolly shouldnt of used an a for different things in that should i

OpenStudy (anonymous):

tan(a) = slope of first line tan(b) = slope of second line |a - b| = angle between them

OpenStudy (amistre64):

\[cos(a)=\frac{m_1.m_2}{|m_1||m_2|}\] thats better :)

OpenStudy (anonymous):

You can use arctan to figure this out. Drawing the lines out and the corresponding triangles should help you visualize the problem.

OpenStudy (amistre64):

the russians method is quite elegant as well

OpenStudy (anonymous):

I just love when someone asks a question and then never responds

OpenStudy (anonymous):

a-b formula is not working I found the slopes but I don't know after that

OpenStudy (anonymous):

one you find the slopes, take the arctan of of each and subtract them